What is equation of motion of lagrage.
1
Expert's answer
2020-12-08T06:29:06-0500

Let`s write lagrange equation of motion:

L = T – V (Kinetic – Potential energies),

Lagrange equation for conservative systems:

"\frac{d}{dt}(\frac{\delta L}{\delta q})-\frac{\delta L}{\delta q}=0".